# Enhancing key rates of QKD protocol by Coincidence Detection

### v1

@@ Abstract
In theory, quantum key distribution (QKD) provides unconditional security; however, its practical implementations are susceptible to exploitable vulnerabilities. This investigation tackles the constraints in practical QKD implementations using weak coherent pulses. We improve on the conventional approach of using decoy pulses by integrating it with the coincidence detection (CD) protocol. Additionally, we introduce an easy-to-implement algorithm to compute asymptotic key rates for the protocol. Furthermore, we have carried out an experimental implementation of the protocol, where we demonstrate that monitoring coincidences in the decoy state protocol leads to enhanced key rates under realistic experimental conditions.
@@ I Introduction
To address the exposed vulnerabilities in the practical implementation of QKD protocols, we can either transition to a more robust protocol or apply security patches for a known attack. Security patching involves monitoring various parameters of our system and assessing potential information leaks. By accounting for imperfections in the source, we can estimate the information leakage caused by pulses containing more than one photon. Security proofs that take into account device imperfections were introduced in Refs. Lutkenhaus2000 ; Gottesman ; Inamori2007 . Additionally, the authors in Ref. Hwang2003 proposed employing decoy states as a defence against the PNS attack. By integrating the work of Gottesman and Hwang2003 , the framework for the decoy-state protocol was built Lo2005 ; Wang2005 ; Ma2005 , which is one of the most widely implemented QKD protocols.
@@
In the present study, we propose combining decoy state protocol with coincidence detection to circumvent a large class of such sophisticated PNS attacks. Following the common convention, we refer to the sender as Alice, the receiver as Bob, and the adversary as Eve. We monitor the coincidences not only for the signal but also for the decoy states. Since Eve does not have the information on whether a signal or decoy is sent, it is not possible for Eve to practically mimic the statistics at the receiver’s end. This is done in addition to the standard decoy state protocol, hence adding an extra layer of security. Therefore, even if Eve manages to accurately reproduce the statistics on average at Bob’s side using most strategies, this will result in a discrepancy in the yield between the signal and the decoy pulses. By incorporating two-fold coincidences, we can achieve higher key rates while enhancing security against sophisticated PNS attacks facilitated by decoy states. The integration of decoy states and coincidence detections helps us to strategically leverage their capabilities to our advantage. We will call this decoy state coincidence detection (DSCD) protocol.
@@ II Protocol
The implementation of the BB84 protocol is susceptible to PNS attacks, where Eve exploits occasional multi-photon pulses generated by Alice’s source. The conventional countermeasure to the PNS attack, namely the decoy state protocol, involves modifications at Alice’s end by sending occasional decoy pulses alongside signal pulses. An alternative approach to combat the PNS attack is to monitor the coincidences at Bob’s end, as suggested by the works of Lim et al. Lim:17 and Biswas et al. biswas_quantum_2022 . The protocol proposed in this paper implements modifications at both Alice’s and Bob’s ends. Decoy states are sent from Alice’s end, and Bob’s end is equipped to monitor coincidence detections, effectively addressing PNS attacks and significantly enhancing key rates. For clarity, the protocol is systematically stated below:
Protocol (Decoy State Coincidence Detection Protocol (DSCD)).
Here, we outline a general decoy state protocol, including coincidences monitoring. Alice transmits a signal with mean photon number $\nu_{0}$ accompanied by $K$ random decoy pulses with mean photon numbers $\nu_{1},\nu_{2},\cdots\nu_{K}$. Given a sequence of $n$ pulses, each labelled with an index $i$, the following steps are repeated for each pulse.
@@
    State preparation (Alice’s lab): Alice generates random numbers $\mathcal{D}_{i}\in\{0,1,\cdots K\}$, $\mathcal{X}_{i}\in\{0,1\}$ and $\mathcal{A}_{i}\in\{0,1\}$. Where,
@@
      $\mathcal{D}_{i}$ determines whether a signal ($\nu_{0}$) or a decoy-state ($\nu_{1},\nu_{2},\cdots\nu_{K}$) is transmitted. If $\mathcal{D}=d$ then mean photon number is $\nu_{d}$
@@
    State measurement (Bob’s lab): Bob lets the pulse pass through a beam splitter, having two ports $\mathfrak{B}^{(0)}$ and $\mathfrak{B}^{(1)}$. At wing, $\mathfrak{B}^{(0)}$ ($\mathfrak{B}^{(1)}$) Bob performs measurement in the Standard (Hadamard) basis. Let $\mathcal{B}_{i}^{(j)}$ denote the measured bit value for the pulse $i$ in port $j$.
@@
      If $\mathcal{B}_{i}^{(j\oplus 1)}=\perp$ and $\mathcal{B}_{i}^{(j)}\in\{0,1\}$: $\mathcal{Y}_{i}=j$ , $\mathcal{B}_{i}=\mathcal{B}_{i}^{(j)}$.
@@
      If $\mathcal{B}_{i}^{(0)},\mathcal{B}_{i}^{(1)}\in\{0,1\}$, then Bob sets $\mathcal{Y}_{i}=x$ and $\mathcal{B}_{i}=\mathcal{B}_{i}^{(x)}$ if $\mathcal{X}_{i}=x$.
@@
    Alice also discloses the decoy states used in each round $(\mathcal{D}_{1},\mathcal{D}_{2},\cdots,\mathcal{D}_{n})$.
@@
    and QBERs (the probability that $\mathcal{A}_{i}\neq\mathcal{B}_{i}$ whenever $\mathcal{X}_{i}=\mathcal{Y}_{i}$ ):
@@
$R\geq\frac{1}{2}\left\{-Q_{\mu}H_{\text{bin}}(E_{\mu})f(E_{\mu})+Q_{1}\left(1-\Phi\left(2e_{1}-1\right)\right)+Q_{2}\left(1-\Phi\left((2e_{2}-1)^{2}\right)\right)\right\}.$ (5)
@@
The standard decoy state protocol usually outperforms the BB84 protocol in key rate due to improved estimation of single-photon statistics ($Q_{1}$ and $e_{1}$). Decoy state statistics aid in obtaining better estimates for both signal and decoy states, resulting in higher key rates. The same argument also applies in our protocol, where having the decoy states, in principle, also allows for better estimates for the single and two-photon statistics $Q_{1},Q_{2},e_{1}$ and $e_{2}$.
Note that, as pointed out in felix_faint_2001 ; biswas_quantum_2022 ; Lim:17 , the PNS attack itself can be detected by simply measuring the coincidences even in the absence of decoy states. However, there are modifications to the PNS attack that cannot be detected simply by monitoring coincidences. One such example is an attack in which Eve only allows for one photon to pass and then sends photons herself to match the ideal coincidence detection. However, such an attack is easily prevented by using decoy states. This is because the adversary does not fundamentally know whether the state sent is a decoy or not; therefore, she will not be able to match the coincidences correctly when Alice and Bob condition on the rounds in which the signal/decoy was sent.
However, the strongest form of the modifications in the PNS attack will still not be detected using our protocol. In this attack, Eve injects a photon only if she stores one. However, we argue that such an attack is not practical for Eve to implement given the current technology. For Eve to execute such an attack, she must first determine that she has stored a photon, generate a replacement photon, and then send it to Bob. This process, especially with current technology, is likely to create a detectable time delay, which Bob can identify. Therefore, ideally, Bob should utilize the time lag information to discard rounds with significant time discrepancies.
@@ III Computing key-rate
$r:=\min$ $\Big{(}Y_{1}\mu\left(1-\Phi\left(2e_{1}-1\right)\right)+Y_{2}\frac{\mu^{2}}{2}\left(1-\Phi\left((2e_{2}-1)^{2}\right)\right)\Big{)}$ (6)
$\mathrm{s.t.}$ $\forall k:\quad Y_{k},e_{k}\in[0,1]$
@@
$\forall d:\quad E_{\nu_{d}}Q_{\nu_{d}}e^{\nu_{d}}=\sum_{k=0}^{\infty}e_{k}Y_{k}\frac{\nu_{d}^{k}}{k!}$
The reason for performing such an optimization is straightforward to understand. The objective function of the optimization problem concerns the key rate $R$ (refer to Eqn (5)). This key rate relies on the $k$ photon yields $Y_{k}$ and $k$ photon error rates $e_{k}$, which are not directly known from experimental statistics. However, what is known are the overall gain $Q_{\mu}$ and the QBER ($E_{\mu}$). The constraints of the optimization problem are related to the observed Gain and QBER for both the signal and decoy states. The fundamental idea is that the experimentally observed QBER and Gain should originate from certain, albeit unknown, fixed values of Yields and error rates.
Given that no assumptions can be made about the Yields $Y_{k}\in[0,1]$ and error rates $e_{k}\in[0,1]$, we acknowledge the possibility of a potential adversary to manipulate them freely. The adversary is, in principle, unrestricted in choosing any values for these parameters, provided that the chosen values align with the observed experimental statistics. This fact is captured using the constraints in the optimization problem, limiting the choice available to the adversary. To account for the strategy that is most advantageous to the adversary within these constraints, we minimize the rate. In the literature, Ma et al. Ma2005 conducted the first comprehensive exploration of lower bounds on the asymptotic key rates for the decoy state protocol. For the decoy state protocol without any coincidence detection, the optimization problem for computing the key rate differs from (6) only in the objective function. Specifically, it does not include contributions from the two-photon terms $Y_{2}$ and $e_{2}$ and is given by:
@@
$Y_{1}\geq Y^{L}_{1}$ ::=:= $\frac{\mu}{\mu\nu-\nu^{2}}\left(Q_{\nu}e^{\nu}-Q_{\mu}e^{\mu}\left(\frac{\nu^{2}}{\mu^{2}}\right)-Y_{0}^{U}\left(\frac{\mu^{2}-\nu^{2}}{\mu^{2}}\right)\right),$ (9)
@@
$R\geq\frac{1}{2}\Big{\{}-Q_{\mu}f(E_{\mu})H_{\mathrm{bin}}(E_{\mu})+Y_{1}^{L}\mu e^{-\mu}\left(1-H_{\mathrm{bin}}(e_{1}^{U})\right)\Big{\}}.$ (11)
An immediate approach to computing lower bounds on the key rate for the protocol with coincidence detection involves employing similar analytic techniques to establish upper and lower bounds, denoted as $Y_{2}^{L}$ and $e_{2}^{U}$, on $Y_{2}$ and $e_{2}$ respectively, that are compatible with the constraints. However, this process is tedious, and tight bounds may not be easily found. While numerical methods can be employed to find such bounds (one approach to obtain such bounds is shown in Appendix D), the results obtained are not sufficiently tight, as demonstrated in (refer to Figure-2(a). In the next section, we introduce a simple-to-implement method to iteratively derive bounds on the key rate that converge to the asymptotic key rate from below.
@@ III.1 Solving the optimization problem
Two primary challenges emerge when addressing optimization problem (6). First, the problem is non-convex due to the non-convex nature of both the constraints and the objective function in (6). General optimization problems are notoriously difficult to solve due to their abstract nature unless they fit specific classes of convex optimization problems. As a result, the optimization problem cannot be solved directly using conventional techniques. Second, optimization problem (6) involves infinitely many free parameters, namely, $Y_{k}$ and $e_{k}$. The presence of these infinitely many parameters adds further complexity to the problem.
The challenge of dealing with infinitely many variables can be addressed straightforwardly by recognizing that the contribution of $Y_{k}$ and $e_{k}$ when $k\gg 1$ is negligible to the sums $\sum_{k}(\nu_{d}^{k}/k!)Y_{k}$ and $\sum_{k}(\nu_{d}^{k}/k!)Y_{k}e_{k}$. Consequently, we relax the constraints of (6) by truncating the infinite sums $\sum_{k}(\nu_{d}^{k}/k!)Y_{k}$ to involve sums over a finite number of variables. This comes, however, with a small penalty that depends on the number of terms kept in the sum. We formally implement this truncation in Appendix C.1.
@@
$\sum_{k\in\{1,2\}}Y_{k}\frac{\mu^{k}}{k!}(1-\Phi((2e_{k}-1)^{k}))\geq\sum_{k\in\{1,2\}}Y_{k}\frac{\mu^{k}}{k!}\cdot(1-\xi^{\max}_{k,i}).$ (12)
@@
All the results lead to the following final result which forms the basis of the algorithm for computing key-rate $r$ (as defined in equation (6)). We state it here informally as:
@@
Assuming a random background with errors occurring equally likely, we explicitly set the background error rate as $e_{0}=1/2$. To account for the impact of finite statistics on the experimental values $Q_{\mu}$, $Q_{\nu}$, $E_{\mu}$, and $E_{\nu}$, it is imperative to consider uncertainties in their measurement. In Appendix E, we modify Theorem 1 to accommodate these statistical uncertainties in the computation of rates. It is important to note that the key rates reported here are key rates in the asymptotic limit. However, this analysis only serves to acknowledge the uncertainty associated with experimentally observed values, and the computed rates are still asymptotic key rates compatible with experimentally obtained statistics with relevant uncertainties in observed values. The computation of finite-round statistics can, in principle, be performed using theoretical tools such as the Generalized Entropy Accumulation Theorem GEAT ; GEAT-security . Such analysis would require a more sophisticated analysis and is reserved for future work.

### v2

@@ Abstract
In theory, quantum key distribution (QKD) provides unconditional security; however, its practical implementations are susceptible to exploitable vulnerabilities. This investigation tackles the constraints in practical QKD implementations using weak coherent pulses. We improve on the conventional approach of using decoy pulses by integrating it with the coincidence detection (CD) protocol. Additionally, we employ an easy-to-implement algorithm to compute asymptotic key rates for the protocol. Furthermore, we have carried out an experimental implementation of the protocol, where we demonstrate that monitoring coincidences in the decoy state protocol leads to enhanced key rates under realistic experimental conditions.
@@ I Introduction
To address the exposed vulnerabilities in the practical implementation of QKD protocols, we can either transition to a more robust protocol or apply security patches for known attacks. Security patching involves monitoring various parameters of our system and assessing potential information leaks. By accounting for imperfections in the source, we can estimate the information leakage caused by pulses containing more than one photon. Security proofs that take into account device imperfections were introduced in Refs. Lutkenhaus2000 ; Gottesman ; Inamori2007 . Additionally, the authors in Ref. Hwang2003 proposed employing decoy states as a defence against the PNS attack. By integrating the work of Gottesman and Hwang2003 , the framework for the decoy-state protocol was built Lo2005 ; Wang2005 ; Ma2005 , which is one of the most widely implemented QKD protocols.
@@
In the present study, we propose combining decoy state protocol with coincidence detection to circumvent a large class of such sophisticated PNS attacks. Following the common convention, we refer to the sender as Alice, the receiver as Bob, and the adversary as Eve. We monitor the coincidences not only for the signal but also for the decoy states. Since Eve does not have the information on whether a signal or decoy is sent, it is not possible for Eve to practically mimic the statistics at the receiver’s end. This is done in addition to the standard decoy state protocol, hence adding an extra layer of security. Therefore, even if Eve manages to accurately reproduce the statistics on average at Bob’s side using most strategies, this will result in a discrepancy in the yield between the signal and the decoy pulses. By incorporating two-fold coincidences, we can achieve higher key rates while enhancing security against sophisticated PNS attacks facilitated by decoy states. The integration of decoy states and coincidence detections helps us to strategically leverage their capabilities to our advantage. We will refer to these extra pulses as entrapped ones to avoid confusion with the decoy state protocol. We will call this the entrapped pulse coincidence detection (EPCD) protocol. This integration of entrapped pulses and coincidence detections allows us to strategically exploit their functionalities.
@@ II Protocol
The implementation of the BB84 protocol with weak coherent pulses is susceptible to PNS attacks, where Eve occasionally exploits multi-photon pulses generated by Alice’s source. The conventional countermeasure to the PNS attack, namely the decoy state protocol, involves modifications at Alice’s end by sending occasional decoy pulses alongside signal pulses. An alternative approach to combat the PNS attack is to monitor the coincidences at Bob’s end, as suggested by the works of Lim et al. Lim:17 and Biswas et al. biswas_quantum_2022 . The protocol proposed in this paper implements modifications at both Alice’s and Bob’s ends, aiming for higher key rates. Entrapped pulses are sent from Alice’s end, and Bob’s end is equipped to monitor coincidence detections, effectively addressing PNS attacks and significantly enhancing key rates. For clarity, the protocol is systematically stated below:
Protocol (Entrapped Pulse Coincidence Detection Protocol (EPCD)).
In this protocol, we present a comprehensive approach to utilizing entrapped pulses, along with monitoring coincidences. Alice sends a signal characterized by an average photon number $\nu_{0}$, along with $K$ entrapped pulses characterized by average photon numbers $\nu_{1},\nu_{2},\cdots,\nu_{K}$. For a sequence of $n$ pulses, each identified by an index $i$, the subsequent actions are iterated for each pulse.
@@
    State preparation (Alice’s lab): Alice generates random numbers $\mathcal{D}_{i}\in\{0,1,\cdots K\}$, $\mathcal{X}_{i}\in\{0,1\}$ and $\mathcal{A}_{i}\in\{0,1\}$ (Fig. 2). Where,
@@
      $\mathcal{D}_{i}$ determines whether a signal ($\nu_{0}$) or an entrapped pulse ($\nu_{1},\nu_{2},\cdots\nu_{K}$) is transmitted. If $\mathcal{D}=d$ then mean photon number is $\nu_{d}$
@@
    Figure 2: Graphical representation of signal and entrapped pulses randomly transmitted by Alice
@@
    State measurement (Bob’s lab): Bob lets the pulse pass through a beam splitter, having two ports $\mathfrak{B}^{(0)}$ and $\mathfrak{B}^{(1)}$. At wing, $\mathfrak{B}^{(0)}$ ($\mathfrak{B}^{(1)}$) Bob performs measurement in the Standard (Hadamard) basis. In our experiment, Standard is the Horizontal/Vertical basis, and Hadamard is the Diagonal/Antidiagonal basis. Let $\mathcal{B}_{i}^{(j)}$ denote the measured bit value for the pulse $i$ in port $j$ (Fig. 3).
    Figure 3: Graphical representation of basis selection and recorded measurements at Bob’s end for $i^{th}$ pulse
@@
      If $\mathcal{B}_{i}^{(j\oplus 1)}=\perp$ and $\mathcal{B}_{i}^{(j)}\in\{0,1\}$: $\mathcal{Y}_{i}=j$ , $\mathcal{B}_{i}=\mathcal{B}_{i}^{(j)}$. This indicates a single detection.
@@
      If $\mathcal{B}_{i}^{(0)},\mathcal{B}_{i}^{(1)}\in\{0,1\}$, then Bob sets $\mathcal{Y}_{i}=x$ and $\mathcal{B}_{i}=\mathcal{B}_{i}^{(x)}$ if $\mathcal{X}_{i}=x$. This indicates two-fold coincidence detection.
@@
    Alice also discloses the entrapped pulses used in each round $(\mathcal{D}_{1},\mathcal{D}_{2},\cdots,\mathcal{D}_{n})$.
@@
    and QBERs (the probability that $\mathcal{A}_{i}\neq\mathcal{B}_{i}$ whenever $\mathcal{X}_{i}=\mathcal{Y}_{i}$, i.e. an incorrect measurement for a correct choice of basis):
@@
$\begin{split}R\geq\frac{1}{2}\biggl{\{}&-Q_{\mu}H_{\text{bin}}(E_{\mu})f(E_{\mu})\\ &+Q_{1}\left(1-\Phi\left(2e_{1}-1\right)\right)\\ &+Q_{2}\left(1-\Phi\left((2e_{2}-1)^{2}\right)\right)\biggr{\}}.\end{split}$ (5)
@@
The standard decoy state protocol usually outperforms the BB84 protocol in key rate due to improved estimation of single-photon statistics ($Q_{1}$ and $e_{1}$). Decoy state statistics aid in obtaining better estimates for both signal and decoy states, resulting in higher key rates. The same argument also applies in our protocol, where having the entrapped pulses, in principle, also allows for better estimates for the single and two-photon statistics $Q_{1},Q_{2},e_{1}$ and $e_{2}$.
Note that, as pointed out in felix_faint_2001 ; biswas_quantum_2022 ; Lim:17 , the PNS attack itself can be detected by simply measuring the coincidences even in the absence of entrapped pulses. However, there are modifications to the PNS attack that cannot be detected simply by monitoring coincidences. One such example is an attack in which Eve only allows for one photon to pass and then sends photons herself to match the ideal coincidence detection. However, such an attack is easily prevented by using entrapped pulses in addition to monitoring coincidences. This is because the adversary cannot distinguish between the entrapped pulses and the signal state; therefore, she will not be able to match the coincidences correctly.
However, the strongest form of the modifications in the PNS attack will still not be detected using our protocol. In this attack, Eve injects a photon only if she stores one. However, we argue that such an attack is not practical for Eve to implement, given the current technology. For Eve to execute such an attack, she must first determine that she has stored a photon, generate a replacement photon, and then send it to Bob. This process, especially with current technology, is likely to create a detectable time delay, which Bob can identify. Therefore, ideally, Bob should utilize the time lag information to discard rounds with significant time discrepancies.
@@ III Computing key-rate
$r:=\min\Big{(}$ $Y_{1}\mu\left(1-\Phi\left(2e_{1}-1\right)\right)$
$+Y_{2}\frac{\mu^{2}}{2}\left(1-\Phi\left((2e_{2}-1)^{2}\right)\right)\Big{)}$
s.t. $\forall k:\quad Y_{k},e_{k}\in[0,1]$
@@
$\forall d:\quad E_{\nu_{d}}Q_{\nu_{d}}e^{\nu_{d}}=\sum_{k=0}^{\infty}e_{k}Y_{k}\frac{\nu_{d}^{k}}{k!}$ (6)
The reason for performing such an optimization is straightforward to understand. The objective function of the optimization problem concerns the key rate $R$ (refer to Eqn (5)). This key rate relies on the $k$ photon yields $Y_{k}$ and $k$ photon error rates $e_{k}$, which are not directly known from experimental statistics. However, what is known are the overall gain $Q_{\mu}$ and the QBER ($E_{\mu}$). The constraints of the optimization problem are related to the observed Gain and QBER for both the signal and entrapped pulses. The fundamental idea is that the experimentally observed QBER and Gain should originate from certain, albeit unknown, fixed values of Yields and error rates.
Given that no assumptions can be made about the Yields $Y_{k}\in[0,1]$ and error rates $e_{k}\in[0,1]$, we acknowledge the possibility of a potential adversary to manipulate them freely. The adversary is, in principle, unrestricted in choosing any values for these parameters, provided that the chosen values align with the observed experimental statistics. This fact is captured using the constraints in the optimization problem, limiting the choice available to the adversary. To account for the strategy that is most advantageous to the adversary within these constraints, we minimize the rate. In the literature, Ma et al. Ma2005 conducted the first comprehensive exploration of lower bounds on the asymptotic key rates for the decoy state protocol. For the decoy state protocol without any coincidence detection, the optimization problem for computing the key rate differs from (III) only in the objective function. Specifically, it does not include contributions from the two-photon terms $Y_{2}$ and $e_{2}$ and is given by:
@@
$Y_{1}\geq Y^{L}_{1}$ ::=:= $\frac{\mu}{\mu\nu-\nu^{2}}\left(Q_{\nu}e^{\nu}-Q_{\mu}e^{\mu}\left(\frac{\nu^{2}}{\mu^{2}}\right)\right.$ (9)
$\left.-Y_{0}^{U}\left(\frac{\mu^{2}-\nu^{2}}{\mu^{2}}\right)\right),$
@@
[…]
$+Y_{1}^{L}\mu e^{-\mu}\left(1-H_{\mathrm{bin}}(e_{1}^{U})\right)\Big{\}}.$ (11)
An immediate approach to computing lower bounds on the key rate for the protocol with coincidence detection involves employing similar analytic techniques to establish upper and lower bounds, denoted as $Y_{2}^{L}$ and $e_{2}^{U}$, on $Y_{2}$ and $e_{2}$ respectively, that are compatible with the constraints. However, this process is tedious, and tight bounds may not be easily found. While numerical methods can be employed to find such bounds (one approach to obtain such bounds is shown in Appendix D), the results obtained are not sufficiently tight, as demonstrated in (refer to Figure-5(a). In the next section, we introduce a simple-to-implement method to iteratively derive bounds on the key rate that converge to the asymptotic key rate from below.
@@ III.1 Solving the optimization problem
Two primary challenges emerge when addressing optimization problem (III). First, the problem is non-convex due to the non-convex nature of both the constraints and the objective function in (III). General optimization problems are notoriously difficult to solve due to their abstract nature unless they fit specific classes of convex optimization problems. As a result, the optimization problem cannot be solved directly using conventional techniques. Second, optimization problem (III) involves infinitely many free parameters, namely, $Y_{k}$ and $e_{k}$. The presence of these infinitely many parameters adds further complexity to the problem.
The challenge of dealing with infinitely many variables can be addressed straightforwardly by recognizing that the contribution of $Y_{k}$ and $e_{k}$ when $k\gg 1$ is negligible to the sums $\sum_{k}(\nu_{d}^{k}/k!)Y_{k}$ and $\sum_{k}(\nu_{d}^{k}/k!)Y_{k}e_{k}$. Consequently, we relax the constraints of (III) by truncating the infinite sums $\sum_{k}(\nu_{d}^{k}/k!)Y_{k}$ to involve sums over a finite number of variables. This comes, however, with a small penalty that depends on the number of terms kept in the sum. We formally implement this truncation in Appendix C.1.
@@
${}\!\!\!\!\!\!\!\!\sum_{k\in\{1,2\}}Y_{k}\frac{\mu^{k}}{k!}(1-\Phi((2e_{k}-1)^{k}))$
$\geq\sum_{k\in\{1,2\}}Y_{k}\frac{\mu^{k}}{k!}\cdot(1-\xi^{\max}_{k,i}).$ (12)
@@
All the results lead to the following final result, which forms the basis of the algorithm for computing key rate $r$ (as defined in equation (III)). We state it here informally as:
@@
Assuming a random background with errors occurring equally likely, we explicitly set the background error rate as $e_{0}=1/2$. To account for the impact of finite statistics on the experimental values $Q_{\mu}$, $Q_{\nu}$, $E_{\mu}$, and $E_{\nu}$, it is imperative to consider uncertainties in their measurement. In Appendix E, we modify Theorem 1 to accommodate these statistical uncertainties in the computation of rates. It is important to note that the key rates reported here are key rates in the asymptotic limit. However, this analysis only serves to acknowledge the uncertainty associated with experimentally observed values, and the computed rates are still asymptotic key rates compatible with experimentally obtained statistics with relevant uncertainties in observed values. The computation of finite-round statistics can, in principle, be performed using theoretical tools such as the Generalized Entropy Accumulation Theorem GEAT ; GEAT-security . Such analysis would require more sophisticated analysis and is reserved for future work.